OpenStudy (anonymous):

Use the quadratic formula to solve x2 + 5x = –2.

OpenStudy (mimi_x3):

\(x^2+5x+2=0\) => \(a=1, b=5, c=2\) Then use the quadratic formula

Parth (parthkohli):

That leaves this for us: \( \color{Black}{\Rightarrow x^2 + 5x + 2 = 0 }\) Quadratic formula is: \( \color{Black}{\Rightarrow x = \Large {-b \pm \sqrt{b^2 - 4ac} \over 2a} }\)

OpenStudy (anonymous):

thankyou